OpenStudy (anonymous):

solve the equation. 2(3x+5)=46 a. 4 b. 5 c. 6 d. 7

OpenStudy (mathmale):

Now check this solution by substituting x=7 into the original equation.

undeadknight26 (undeadknight26):

oh its C...

OpenStudy (mathmale):

Are you concluding that x = 7 is correct or incorrect? Please be specific.

undeadknight26 (undeadknight26):

Its incorrect -_- because the right answer is C. or x = 6...

OpenStudy (khaakenajaf):

step by step is best way to solve qestions

OpenStudy (khaakenajaf):

2(3x+5)=46 divide both sides by 2 3x+5=23 3x=23-5 3x=18 x=6 answer is C verify the answer 2(3*(6)+5)=46 2(18+5)=46 2*23=46 46=46 verified